OpenStudy (anonymous):

Is the function x^(1/3) differentiable at 0? why or why not?

myininaya (myininaya):

try to find the derivative function of x^(1/3)

OpenStudy (p0sitr0n):

it is continuous, so yes

myininaya (myininaya):

but is the function smooth there?

myininaya (myininaya):

The function needs to be continuous and smooth at x=a to be differentiatiable at x=a

OpenStudy (anonymous):

derivative of x^1/3= 1/(3x^(2/3))

OpenStudy (anonymous):

Now how do I prove that it's differentiable?

OpenStudy (anonymous):

if it is continuous so yes? this is much easier than that take the derivative if you can evaluate it at zero, then yes if you cannot, then no

OpenStudy (anonymous):

smooth... like gary coleman

myininaya (myininaya):

lol

ganeshie8 (ganeshie8):

you may use limit definition to make your proof a bit more lengthy

myininaya (myininaya):

@aawowaa you can't divide by 0

myininaya (myininaya):

But if you are looking for something more formal there is that thingy that @ganeshie8 mentioned

ganeshie8 (ganeshie8):

derivative = 1/0 at a point tells us that the tangent line is vertical

OpenStudy (anonymous):

So the function is not differentiable... right?

ganeshie8 (ganeshie8):

derivative of x^1/3= `1/(3x^(2/3))` show that this value is not defined at x = 0

OpenStudy (anonymous):

1/(3(0)^(2/3))=1/0 so it's undefined

OpenStudy (aum):

Correct. Differentiable means the derivative exists. At x = 0, the derivative is undefined. Therefore, at x = 0, the function is not differentiable.